OpenStudy (anonymous):

Given that sin θ = sqrt(3)/7 and tan θ < 0, find sec θ and tan θ.

OpenStudy (y2o2):

θ is in the second quad. so you can use sin^-1 function or you can use Pythagoras theorem and SOH CAH TOA

OpenStudy (anonymous):

ok so x^2 + y^2 = r^2 x^2 + sqrt(3)^2 = 7^2 x^2 + 3 = 49 x^2 = 46 x = sqrt(46) right?

OpenStudy (y2o2):

yes you got the adjacent right. so now you can get the cos and the tan but remember that both of them will be -ve

OpenStudy (anonymous):

So Sec θ = - 7sqrt(46)/46 Tan θ = - sqrt(138)/46

OpenStudy (y2o2):

Yes, that's right

OpenStudy (anonymous):

cool, thanks. I wasn't sure if I could reduce it further.